OpenStudy (anonymous):

If the mass of an object is 52kg, what does it weigh?

OpenStudy (anonymous):

weight is a force = m * g where g = acceleration due to gravity (ABOUT 9.8 M/S^2) so here weight = 52 * 9.8 newtons

OpenStudy (anonymous):

Thank you.

OpenStudy (asnaseer):

You might also find this useful: http://www.colorado.edu/physics/2000/periodic_table/mass.html